# Towards Interactive and Learnable Cooperative Driving Automation: a Large Language Model-Driven Decision-Making Framework

### v2

@@ IV-D2 Safety Evaluation
We have conducted a comprehensive safety analysis of various methods, utilizing the Post-Encroachment Time (PET) metric, a widely recognized safety parameter in traffic engineering. This metric effectively quantifies the safety and interaction intensity of vehicles in intricate traffic scenarios. According to the results presented in Fig. 9 and Table. I, both the iDFST approach and CoDrivingLLM exhibit commendable performance, with mean PET values of 15.1 seconds and 10.3 seconds, respectively. Conversely, MADQN and the Cooperative method demonstrate PET values of 5.7 seconds and 3.7 seconds, respectively, indicating a heightened risk in decision-making interactions within the traffic environment. In addition, a lower PET value signifies an elevated safety hazard. Normally, a PET value less than 1.5s signifies severe conflicts that should be avoided. According to Fig. 9, The PET distribution of the proposed CoDrivingLLM is predominantly above 1.5s. Although the PET distribution for MADQN shows no values below 1.5s as well, considering that its success rate is only 20%, there are a large number of collisions. Therefore, CoDrivingLLM performs best in terms of safety.
@@ IV-D3 Efficiency Evaluation
We also evaluate the efficiency of all methods using travel velocity. As shown in Table. I, MADQN achieves the highest average speed at 6.1 m/s during the tests. However, this efficiency comes at the cost of significantly compromising safety and overall performance. The iDFST method performs well in terms of safety and success rate, it exhibits poor driving efficiency, with excessive low-speed driving resulting in an average speed of only 4.1 m/s. Although the rule-based approach performs well in terms of safety, its low efficiency has consistently been a common issue. In contrast, CoDrivingLLM strikes a balance between safety and efficiency, maintaining a high average speed of 5.7 m/s while ensuring safety.
@@ V Conclusion
In the future, we aim to broaden the application of this framework to encompass a wider range of scenarios. We plan to undertake large-scale cooperative decision-making experiments involving CAVs, and we aspire to deploy this framework on actual experimental vehicles in real-world settings. Additionally, we intend to refine the memory module and evaluate various LLM base models to bolster the practical applicability and performance of our approach.

### v3

@@ IV-D2 Safety Evaluation
Furthermore, we conducted a more in-depth analysis of the performance of different methods in terms of safety and efficiency. In this analysis, only successful cases were considered when calculating the relevant metrics. We have conducted a comprehensive safety analysis of various methods, utilizing the Post-Encroachment Time (PET) metric, a widely recognized safety parameter in traffic engineering. This metric effectively quantifies the safety and interaction intensity of vehicles in intricate traffic scenarios. According to the results presented in Fig. 9 and Table. I, the proposed CoDrivingLLM, FCFS, iDFST, and MCTS all show relatively high average PET values, indicating good stability in terms of safety. In contrast, Cooperative Game, MADQN, and Dilu exhibit average PET values of less than 6 seconds, suggesting that these methods involve a significant number of interactions with lower PET, which increases the risk of driving hazards.
Additionally, the minimum PET reflects the most dangerous interactions in collision-free cases, highlighting the critical moments when the risk of conflict is highest. Normally, a PET value less than 1.5s signifies severe conflicts that should be avoided [43]. According to Fig. 9, The PET distribution of the proposed CoDrivingLLM is predominantly above 1.5s. Although the PET distribution for MADQN shows no values below 1.5s as well, considering that its success rate is only 20%, there are a large number of collisions. Therefore, CoDrivingLLM performs best in terms of safety.
@@ IV-D3 Efficiency Evaluation
We also evaluated the efficiency of all methods using travel velocity. As shown in Table. I, MADQN achieves the highest average speed at 6.1 m/s during the tests. However, this efficiency comes at the cost of significantly compromising safety and overall performance. The proposed CoDrivingLLM achieves an average travel velocity of 5.7 m/s while ensuring safety, making it the most efficient method after excluding MADQN, which has a success rate of only 20%.
In addition, MCTS and Cooperative game demonstrate mid-range efficiency, partly due to their optimization objective, which includes minimizing delays. On the other hand, FCFS and iDFST, both rule-based methods, exhibit the lowest efficiency. Notably, while FCFS achieves the same overall success rate as CoDrivingLLM, its average travel velocity is only 2.7 m/s. Lastly, since Dilu is a single-vehicle decision-making method, it lacks information sharing and negotiation. As a result, when applied to intersection scenarios, it leads to frequent deadlocks, resulting in poor efficiency.
We further compared the success rate of the proposed CoDrivingLLM with single-vehicle decision-making methods, available at here  [^3] https://github.com/FanGShiYuu/CoDrivingLLM/blob/master/videos&data/Comparision.md. However, given that vehicle cooperation naturally provides an informational advantage through state and intent sharing, it is expected that CoDrivingLLM outperforms common single-vehicle decision-making methods. Therefore, we do not elaborate on this further.
@@ V Conclusion
In the future, we will continue to explore the potential applications of CoDriving in other scenarios, including its ability to handle heterogeneous background traffic. Additionally, we plan to conduct field tests to further validate its feasibility in more realistic interaction environments. Additionally, we intend to refine the memory module and evaluate various LLM base models to bolster the practical applicability and performance of our approach.